Fermi National Laboratory
 

Accelerator Update for May 21, 2007 - May 23, 2007

 

Monday May 21

The day shift began with Operations monitoring the Tevatron (TeV) store 5450 (48.26E30) and the Recycler stash (162.2E10), with the Antiproton Source (Pbar) stacking (42.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE experiments taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beam.

At 8:47 AM, a Linac RF station (LRF3) tripped off. It reset.

At 9:03 AM, LRF3 tripped off and would not reset. With the help of an expert the station was reset by 9:10 AM.

RF experts reported at 9:27 AM that the MI RF LCW system had a small leak some where around the MIRF14 Ė 16 cavities in the tunnel. They could see water on the floor via the CCTV link. They would like to access and investigate further.

At 9:46 AM, LRF3 tripped off. Experts reset the station.

Operations terminated store 5450 at 11:14 AM.

Operations had CDF and D0 prepared for an access by 11:23 AM.

TeV experts reported at 11:41 AM that loop controller change had been completed for the TeV C2 power supply and that it was back online.

A MI expert began a multibatch slipstacking study at 12:05 PM.

At 12:22 PM, Operations reported problems with the I- Source, a drop in output. It only lasted for a minute and then it resumed normal operations.

Switchyard (SY) experts began a study period at 12:27 PM.

CDF came out of access at 1:15 PM.

The MI study ended at 1:20 PM. Also, Preaccelerator experts accessed the I- Source dome to check on the extractor power supply.

The Preaccelerator experts switched from the I- Source to the H- Source at 2:02 PM.

Operations put the TeV through a dry squeeze at 2:05 PM.

Linac resumed sending beam at 2:14 PM.

At 2:43 PM, Operations reported that the TeV aborted at sectors E4 and D0. Experts investigated and found that beam was coming from MI into the P1 transfer line at a bad angle. MI and TeV experts continued their investigations.

At 3:41 PM, Operations reported that they had been having problems injecting protons into the TeV. Also, Pbar had been having problems stacking. The search to link these problems narrowed to the MI-52 extraction kicker magnet. Operations contacted an expert.

Operations reported at 4:26 PM that the Recycler had stashed antiprotons twice during the day shift.

An MI expert reported at 4:32 PM that the MI-52 kicker was putting out the proper current.

Operations put the TeV through a wet squeeze at 4:32 PM.

Tuesday May 22

The midnight shift began with Operations monitoring store 5453 (89.37E30) and the stash (72.1E10), with Pbar stacking (42.7E10), and with NuMI and MiniBooNE operating normally.

At 4:05 AM, Operations reported that many booster devices faulted. Operators discovered that a Booster CAMAC crate power supply (#83) had failed.

At 4:17 AM, the Pelletron VME crate hung up and would not reboot remotely. It will require an access into the Pelletron and a local reboot.

Operations reported at 4:22 AM that operators replaced CAMAC crate #83ís power supply.

Operations reported at 7:53 AM that a Pbar expert moved the antiproton target to a new position.

SY experts began a study at 8:03 AM.

Operations reported at 8:05 AM that the Recycler had stashed three times during the midnight shift.

At 8:38 AM, a Pbar expert began some target scans. This caused a slight drop in the stacking rate.

Controls experts reported at 9:26 AM that they had reset the Pelletron VME.

Operations reported at 11:15 AM that the Recycler experts had completed a compound bucket study.

Operations terminated store 5453 at 2:59 PM.

Operations terminated store 5453 at 2:59 PM.

Operators prepared the Meson enclosures M01 through MC6 for access at 3:11 PM.

Operations put the TeV into shot setup at 3:35 PM.

Operations reported at 4:01 PM that the Recycler had stashed three times during the day shift.

At 4:38 PM, Operations reported that the TeV quenched at sectors F1 & 2 during the proton injection. The sequencer didnít increment a device (V:NXBNCH) and that caused the injection of protons on top of protons, which caused the quench.

Operations put the TeV through a dry squeeze at 5:58 PM.

Operations restarted the TeV shot setup at 6:23 PM.

Operations established store 5455 at 8:20 PM with an initial luminosity of 200.92E30.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Wednesday May 23

The midnight shift began with Operations monitoring store 5455 (117.68E30) and the stash (98.3E10), with Pbar stacking (38.6E10), and with NuMI and MiniBooNE taking beam.

At 3:40 AM, BRF9 & 10 tripped off due to a vacuum high-pressure indication. They both reset.

At 6:57 AM, the Linac Buncher tripped off for the first of two times in about 30 minutes. It reset.

Operations reported at 8:03 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator may allow studies that have little impact on stacking and stashing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/23/2007    email Fermilab